OpenStudy (anonymous):

If f(x) = 1/2 x + 3 and g(x) = −4x − 1, find g[f(4)]

OpenStudy (anonymous):

First evaluate f(4) by plugging in 4 for x in f(x): f(4)= 0.5(4) + 3 = 5 Then, use the value you just got, and plug that into the x in g(x): g[f(4)]= -4(5) - 1 = -21 So the correct answer is -21. Note: This is the biology area of OpenStudy... maybe post your question in the algebra area?